OpenStudy (anonymous):

under the QUADRATIC EQUATION(Roots) WHat number is less twice its reciprocal gives 23/5?

OpenStudy (saifoo.khan):

x - 2/x = 23/5

OpenStudy (saifoo.khan):

x^2 -2 = 23x/5

OpenStudy (anonymous):

-2/5 or 5

OpenStudy (saifoo.khan):

x^2 -23x/5 -2 = 0 Now solve the quadratic.

OpenStudy (aravindg):

-2/5 or 5

OpenStudy (saifoo.khan):

and they are right with the roots.